# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-A Survey on WF attacks
Naïve Bayes  [8]
[66]
 [37]
 [58]
SVM  [7, 35]
 [37]
[4]
k-NN  [9]
 [33]
 [35]
 [37]
 [79]
 [80]
Decision Tree  [63]
Random Forest  [36]
Hoeffding Tree  [64]
@@
CNN  [38, 118, 96, 119, 115, 120]
 [12, 39, 40, 33]
 [94, 117, 13]
 [121, 43, 122, 42]
 [123]
 [32]
 [124, 93, 125, 126, 127, 128, 109, 129]
 [130]
 [29]
LSTM  [38]
 [94]
SADE  [38]
 [94]
 [10]
Few-shot Learning  [131, 38, 132]
 [127, 133, 134, 135, 136, 137]
 [95, 138]
 [139, 140]
 [76]
DNN  [111]
 [141]
GAN  [142]
 [143]
Transformer  [43]
 [30]
 [70]
GNN  [31]
 [75]
EDF-S and EDF-P  [144]
LRCT  [44]
Enhanced Technique  [145]
NetCLR  [71]
UNetFormer  [146]
CIL  [147]
@@
Panchenko et al. introduced the first WF attack that accurately classified web pages on Tor in 2009. Since then, various techniques have been developed, broadly categorized into single-tab and multi-tab fingerprinting based on user browsing behavior [148, 149]. Multi-tab WF represents more realistic usage patterns, where users access multiple websites simultaneously in different tabs. However, it introduces added complexity, as traffic from multiple sites is interleaved, making it more difficult to distinguish and identify individual websites. To help researchers choose appropriate models for WF attacks, this paper classifies existing WF attack methods into three categories: traditional ML-based, DL-based, and non-machine learning-based approaches. By incorporating traffic-level characteristics, such as direction, timing, and burst patterns, the proposed classification framework offers a more practical and application-oriented method for selecting suitable models. A summary of the current WF attack techniques and their applied features is presented in TABLE V.
@@
Previous WF attacks on Tor have often relied on self-collected open-world datasets, which vary widely in terms of scale and environmental conditions, such as network settings, browser configurations, and webpage selections [35, 11, 94]. This inconsistency makes it difficult to compare the effectiveness of different WF methods directly. To address misleading claims of high recall or true positive rate (TPR) and low false positive rate (FPR) in existing WF attacks, Wang [150] proposed enhancing classifier performance through the use of precision optimizers based on confidence scoring, distance metrics, and ensemble learning. These optimizers prioritize precision over recall, which is better suited for open-world website fingerprinting (OWF), especially in large-scale surveillance settings where the base rate fallacy, the overestimation of performance caused by the rarity of target websites, can distort evaluation. Further, in 2020, Wang [114] highlighted that prior work often failed to properly account for base rates when calculating precision. To resolve this, the author introduced r-precision, a metric that explicitly incorporates the rarity of sensitive pages, offering a more realistic assessment of attack performance in open-world settings. To support more reliable and reproducible evaluations of WF attacks, this survey provides a comprehensive classification of WF models, dataset scales, and accuracy metrics, aiming to bridge the methodological gaps in prior research.
@@ V-A1 Traditional ML-based WF Attack Models
Traditional ML-based WF attacks represent some of the earliest and most extensively studied approaches in the field of traffic analysis [8, 66, 58, 7, 35]. These techniques rely on classical supervised learning algorithms to classify encrypted traffic traces based on statistical features extracted from network flows. Despite the encryption and anonymity provided by Tor, these ML models can often infer which websites users visit by analyzing observable traffic patterns such as packet sizes, directions, timings, and burst behaviors.
In a typical traditional ML-based WF attack, the attacker first collects labeled traffic traces from a set of monitored websites under controlled conditions. These traces extract a range of hand-crafted features, such as total packet count, inter-arrival times, or traffic burst lengths. These features are then used to train classifiers like SVMs [35, 37, 7], K-NN [9, 63, 37, 33], Random Forests [36], or Naïve Bayes [8, 66, 37], which learn to distinguish between traffic patterns corresponding to different websites. TABLE VI summarizes the existing traditional ML-based WF attacks, reporting their performance in both closed-world and open-world settings using practical evaluation datasets. If the dataset column is marked as none, it indicates that the authors used a self-collected dataset rather than a publicly available one. For other columns, none indicates that the corresponding information was not reported or could not be found.
@@ 90.00 NA NA
k-fingerprinting [36] 2016 Random Forest Wang14 [9] 550 91.50 6000 NA
AdaWFPA [64] 2019 Hoeffding Tree Walkie-Talkie [45] $\times100$ 99.60 $\times100$ 99.40
@@
Naïve Bayes-based WF Attacks: Naïve Bayes is one of the earliest and simplest machine learning algorithms applied to WF attacks. It is a probabilistic classifier based on Bayes’ theorem with the assumption of feature independence, which makes it computationally efficient and easy to implement. In the context of WF, Naïve Bayes classifiers are trained on hand-crafted features extracted from encrypted traffic traces, such as total number of packets, packet size distributions, inter-arrival times, and traffic bursts [37]. These features are used to compute the likelihood of a traffic trace belonging to a particular website, and the classifier selects the website with the highest posterior probability. In 2006, Liberatore and Levine [66] first introduced packet lengths as a powerful WF feature for WF attacks. This paper first applied the Naïve Bayes classifier to map each packet sequence with unique packet lengths. Then, the authors used the Jaccard coefficient as a measurement of the distance between two packet sequences and achieved 96.65% accuracy on the OpenSSH dataset. In 2009, Herrmann et al. [8] performed the WF attack on the Tor network for the first time, which was based on a Naïve Bayes classifier with features from traffic flow direction and the normalized frequency distribution of observable IP packet size. In 2012, Dyer et al. [58] introduced a variable n-gram classifier that employs Naïve Bayes as its underlying machine learning algorithm. Their study found that coarse-grained features (Total transmission time, total per-direction bandwidth, and burst bandwidth) were particularly effective in enabling classifiers to distinguish traffic, even when it was protected by packet padding defenses. In summary, Naïve Bayes was shown to be effective in early WF studies despite its simplicity. However, the NB model assumes that all features are independent, which is rarely true in traffic data. Moreover, Naïve Bayes tends to perform poorly in open-world settings, where distinguishing between monitored and non-monitored sites is significantly more challenging.
@@
SVM-based WF Attacks: SVMs are another widely used method in early WF attacks due to their strong performance in high-dimensional classification tasks and their robustness to overfitting. SVMs are particularly attractive for WF because they perform well with limited training data and effectively handle high-dimensional feature spaces. In 2011, Panchenko et al. [4] were the first to apply SVM to WF attacks by defining features based solely on traffic volume, timing, and direction. Their approach significantly improved Tor website classification accuracy from 3% to 55%, and achieved a 73% recall rate in an open-world setting. Their approach is also the first successful WF attack under realistic open-world conditions. In 2012, Cai et al. [7] introduced an optimal string alignment distance (OSAD) to measure the website fingerprinting distance based on packet size for TCP packet traces. Then, the authors applied SVM to classify and recognize websites and web pages with 90% accuracy. Based on Cai’s work, Wang and Goldberg [35] first introduced a method for representing Tor traffic as a one-dimensional array of size $1\times L$], where each element indicates the direction of a Tor cell: +1 for outgoing and -1 for incoming packets. To standardize input lengths across all traffic traces, sequences were padded with zeros to a fixed length, commonly LL=5000. This 1-D array representation has become a standard input format for many WF attack models. Despite their effectiveness, SVM-based WF attacks also face several challenges: (1) The accuracy of SVM-based WF attacks depends heavily on the selection of features. If the wrong features are chosen, or if features are noisy, the model may fail to classify traffic correctly; (2) Performance can degrade in open-world settings due to the large and dynamic website changes; (3) The model may struggle with scalability as the dataset grows.
@@
K-NN based WF Attacks: K-NN is a non-parametric, lazy learning algorithm used for classification and regression tasks. It works by finding the most similar instances (neighbors) to a given input sample and then assigning the class based on a majority vote from these neighbors. In the case of website fingerprinting, K-NN classifies network traffic based on the similarity of the features (such as packet sizes or packet ordering patterns) to the features of known traffic traces [37]. In 2013, Wang et al. [35] improved the K-NN algorithm by introducing time-based splitting and classification-based splitting to isolate traffic corresponding to single or multiple web pages. Their experiments demonstrated that time-based splitting could achieve up to 97.00% TPR for detecting single-page sequences, while classification-based methods further improved accuracy in cases of overlapping traffic. Later, Wang et al. [9] applied the K-NN classifier to a prominent feature set with weight adjustments in a realistic open-world scenario. Their results showed an 85.00% TPR with a FPR of 0.60%, outperforming previous work, which achieved 83.00% TPR with a FPR of 6.00%.
@@
Previous studies [35, 9, 37] often neglect hyperlink transition information, as it introduces additional “noise” that complicates the classification of the original webpage. To address this issue, Zhuo et al. [79] proposed a WFP technique that integrates PHMM with K-NN, where PHMM is traditionally applied in bioinformatics for DNA sequence analysis. This method accounts for hyperlink transitions, making it more realistic and effective. The technique models websites by considering different webpages and their transitions, allowing for the identification of a sequence of page loads belonging to the same website. Many WF attacks assume websites do not change over time. In reality, websites are highly dynamic, leading to significant intra-class variability that reduces classification accuracy. Cui et al. [80] revisited key assumptions underlying WF attacks, which allow adversaries to infer users’ web activity despite encryption. By incorporating Markov models to capture temporal and probabilistic properties of web traffic, the paper highlights that WF attacks can remain effective under more realistic conditions. To distinguish between single-page and two-page traces, the authors employed a K-NN binary classifier. Their results show that the method can identify whether a trace contains one or two pages with an accuracy of 97.00%. Traditional K-NN classifiers typically rely on the labels of the nearest samples for classification, but the diversity of multi-tab traffic can lead to sample drift, degrading performance. To address this challenge, Zhao et al. [33] integrated proxy-based K-NN and sample-based K-NN to enhance multi-tab webpage identification. Their experiments showed that this hybrid model improved the multi-label recall metric by 88.60% compared to the state-of-the-art attacks.
@@
Tree-based WF Attacks: Some researchers have explored Tree-based models to improve the accuracy of WF attacks. Tree-based models use decision trees [63] or ensembles of trees [62, 65] to learn from features extracted from network traffic. These models recursively split the data based on feature thresholds, aiming to maximize the separation of traffic associated with different websites. In 2015, Kwon et al. [63] used C4.5 and CART decision trees to classify Tor hidden service circuits. The authors demonstrate that even without compromising the Tor network or engaging in active traffic manipulation, it is possible to infer which hidden service a Tor client is accessing. By leveraging decision trees and WF, adversaries can deanonymize users and services with high accuracy (>>99.00%). To address the challenge of a large amount of noisy data, Hayes et al. [36] introduced k-fingerprinting, which leverages random decision forests to classify encrypted or anonymized web traffic. The method transforms packet sequences into distinctive ”fingerprints” by mapping them to the leaf node identifiers produced by a trained random forest. It achieves impressive results, with TPR of up to 85% and FPR as low as 0.02% when identifying monitored hidden services.
@@
The accuracy of WF attack models drops drastically when tested on captured traffic traces of websites due to concept drift. To address the challenges posed by concept drift and the high cost of retraining, Attarian et al. [64] introduced AdaWFPA, an adaptive online WF attack that overcomes the limitations of static models, whose accuracy degrades over time as website traffic patterns change. AdaWFPA leverages adaptive stream mining algorithms, Adaptive Hoeffding Tree (AHT), and Adaptive Hoeffding Option Tree (AHOT), to dynamically update its classifier in real-time, ensuring robust performance against evolving website characteristics. The method is evaluated on Tor traffic datasets, demonstrating superior accuracy, precision, and recall compared to state-of-the-art static attacks, even when tested against defenses like Walkie_Talkie [45].

### v2

@@ V-A Survey on WF attacks
Naïve Bayes  [herrmann2009website]
[liberatore2006inferring]
 [wang2016realistically]
 [dyer2012peek]
SVM  [cai2012touching, wang2013improved]
 [wang2016realistically]
[panchenko2011website]
k-NN  [wang2014effective]
 [zhao2024towards]
 [wang2013improved]
 [wang2016realistically]
 [zhuo2017website]
 [cui2019revisiting]
Decision Tree  [kwon2015circuit]
Random Forest  [hayes2016k]
Hoeffding Tree  [attarian2019adawfpa]
@@
CNN  [hong2024website, shen2023subverting, bhat2018var, yuan2025mw3f, gong2024wfcat, cui2020more]
 [rahman2019tik, shi2025multiscale, zou2024relation, zhao2024towards]
 [rimmer2017automated, guan2021bapm, yin2024traces]
 [wang20202ch, jin2023transformer, pan2023practical, xu2025apwf]
 [qian2024enhancing]
 [deng2024robust]
 [wang2022snwf, sirinam2018deep, pulls2020website, liu2024fingermamba, chen2022few, chen2022srp, deng2023robust, chen2022tform]
 [tian2023tiny]
 [deng2025countmamba]
LSTM  [hong2024website]
 [rimmer2017automated]
SADE  [hong2024website]
 [rimmer2017automated]
 [abe2016fingerprinting]
Few-shot Learning  [zou2022efficient, hong2024website, meng2025beyond]
 [chen2022few, chen2021few2, lyu2024tfan, zhou2024joint, wang2024trafficsiam, tan2024adaptability]
 [sirinam2019triplet, wang2024wf3a]
 [xie2024contrastive, li2025cross]
 [zhu2025wf]
DNN  [oh2017p]
 [singh2023first]
GAN  [oh2021gandalf]
 [xie2023multi]
Transformer  [jin2023transformer]
 [mathews2024laserbeak]
 [zhou2023wf]
GNN  [gao2024multi]
 [tan2024inter]
EDF-S and EDF-P  [cui2023edf]
LRCT  [ding2024multi]
Enhanced Technique  [mei2025high]
NetCLR  [bahramali2023realistic]
UNetFormer  [yang2025dtpn]
CIL  [yuan2025class]
@@
Panchenko et al. introduced the first WF attack that accurately classified web pages on Tor in 2009. Since then, various techniques have been developed, broadly categorized into single-tab and multi-tab fingerprinting based on user browsing behavior [holland2021new, divakaran2025traffic]. Multi-tab WF represents more realistic usage patterns, where users access multiple websites simultaneously in different tabs. However, it introduces added complexity, as traffic from multiple sites is interleaved, making it more difficult to distinguish and identify individual websites. To help researchers choose appropriate models for WF attacks, this paper classifies existing WF attack methods into three categories: traditional ML-based, DL-based, and non-machine learning-based approaches. By incorporating traffic-level characteristics, such as direction, timing, and burst patterns, the proposed classification framework offers a more practical and application-oriented method for selecting suitable models. A summary of the current WF attack techniques and their applied features is presented in TABLE V.
@@
Previous WF attacks on Tor have often relied on self-collected open-world datasets, which vary widely in terms of scale and environmental conditions, such as network settings, browser configurations, and webpage selections [wang2013improved, panchenko2016website, rimmer2017automated]. This inconsistency makes it difficult to compare the effectiveness of different WF methods directly. To address misleading claims of high recall or true positive rate (TPR) and low false positive rate (FPR) in existing WF attacks, Wang [wang2018optimizing] proposed enhancing classifier performance through the use of precision optimizers based on confidence scoring, distance metrics, and ensemble learning. These optimizers prioritize precision over recall, which is better suited for open-world website fingerprinting (OWF), especially in large-scale surveillance settings where the base rate fallacy, the overestimation of performance caused by the rarity of target websites, can distort evaluation. Further, in 2020, Wang [wang2020high] highlighted that prior work often failed to properly account for base rates when calculating precision. To resolve this, the author introduced r-precision, a metric that explicitly incorporates the rarity of sensitive pages, offering a more realistic assessment of attack performance in open-world settings. To support more reliable and reproducible evaluations of WF attacks, this survey provides a comprehensive classification of WF models, dataset scales, and accuracy metrics, aiming to bridge the methodological gaps in prior research.
@@ V-A1 Traditional ML-based WF Attack Models
Traditional ML-based WF attacks represent some of the earliest and most extensively studied approaches in the field of traffic analysis [herrmann2009website, liberatore2006inferring, dyer2012peek, cai2012touching, wang2013improved]. These techniques rely on classical supervised learning algorithms to classify encrypted traffic traces based on statistical features extracted from network flows. Despite the encryption and anonymity provided by Tor, these ML models can often infer which websites users visit by analyzing observable traffic patterns such as packet sizes, directions, timings, and burst behaviors.
In a typical traditional ML-based WF attack, the attacker first collects labeled traffic traces from a set of monitored websites under controlled conditions. These traces extract a range of hand-crafted features, such as total packet count, inter-arrival times, or traffic burst lengths. These features are then used to train classifiers like SVMs [wang2013improved, wang2016realistically, cai2012touching], K-NN [wang2014effective, kwon2015circuit, wang2016realistically, zhao2024towards], Random Forests [hayes2016k], or Naïve Bayes [herrmann2009website, liberatore2006inferring, wang2016realistically], which learn to distinguish between traffic patterns corresponding to different websites. TABLE VI summarizes the existing traditional ML-based WF attacks, reporting their performance in both closed-world and open-world settings using practical evaluation datasets. If the dataset column is marked as none, it indicates that the authors used a self-collected dataset rather than a publicly available one. For other columns, none indicates that the corresponding information was not reported or could not be found.
@@ 90.00 NA NA
k-fingerprinting [hayes2016k] 2016 Random Forest Wang14 [wang2014effective] 550 91.50 6000 NA
AdaWFPA [attarian2019adawfpa] 2019 Hoeffding Tree Walkie-Talkie [wang2017walkie] $\times100$ 99.60 $\times100$ 99.40
@@
Naïve Bayes-based WF Attacks: Naïve Bayes is one of the earliest and simplest machine learning algorithms applied to WF attacks. It is a probabilistic classifier based on Bayes’ theorem with the assumption of feature independence, which makes it computationally efficient and easy to implement. In the context of WF, Naïve Bayes classifiers are trained on hand-crafted features extracted from encrypted traffic traces, such as total number of packets, packet size distributions, inter-arrival times, and traffic bursts [wang2016realistically]. These features are used to compute the likelihood of a traffic trace belonging to a particular website, and the classifier selects the website with the highest posterior probability. In 2006, Liberatore and Levine [liberatore2006inferring] first introduced packet lengths as a powerful WF feature for WF attacks. This paper first applied the Naïve Bayes classifier to map each packet sequence with unique packet lengths. Then, the authors used the Jaccard coefficient as a measurement of the distance between two packet sequences and achieved 96.65% accuracy on the OpenSSH dataset. In 2009, Herrmann et al. [herrmann2009website] performed the WF attack on the Tor network for the first time, which was based on a Naïve Bayes classifier with features from traffic flow direction and the normalized frequency distribution of observable IP packet size. In 2012, Dyer et al. [dyer2012peek] introduced a variable n-gram classifier that employs Naïve Bayes as its underlying machine learning algorithm. Their study found that coarse-grained features (Total transmission time, total per-direction bandwidth, and burst bandwidth) were particularly effective in enabling classifiers to distinguish traffic, even when it was protected by packet padding defenses. In summary, Naïve Bayes was shown to be effective in early WF studies despite its simplicity. However, the NB model assumes that all features are independent, which is rarely true in traffic data. Moreover, Naïve Bayes tends to perform poorly in open-world settings, where distinguishing between monitored and non-monitored sites is significantly more challenging.
@@
SVM-based WF Attacks: SVMs are another widely used method in early WF attacks due to their strong performance in high-dimensional classification tasks and their robustness to overfitting. SVMs are particularly attractive for WF because they perform well with limited training data and effectively handle high-dimensional feature spaces. In 2011, Panchenko et al. [panchenko2011website] were the first to apply SVM to WF attacks by defining features based solely on traffic volume, timing, and direction. Their approach significantly improved Tor website classification accuracy from 3% to 55%, and achieved a 73% recall rate in an open-world setting. Their approach is also the first successful WF attack under realistic open-world conditions. In 2012, Cai et al. [cai2012touching] introduced an optimal string alignment distance (OSAD) to measure the website fingerprinting distance based on packet size for TCP packet traces. Then, the authors applied SVM to classify and recognize websites and web pages with 90% accuracy. Based on Cai’s work, Wang and Goldberg [wang2013improved] first introduced a method for representing Tor traffic as a one-dimensional array of size $1\times L$], where each element indicates the direction of a Tor cell: +1 for outgoing and -1 for incoming packets. To standardize input lengths across all traffic traces, sequences were padded with zeros to a fixed length, commonly LL=5000. This 1-D array representation has become a standard input format for many WF attack models. Despite their effectiveness, SVM-based WF attacks also face several challenges: (1) The accuracy of SVM-based WF attacks depends heavily on the selection of features. If the wrong features are chosen, or if features are noisy, the model may fail to classify traffic correctly; (2) Performance can degrade in open-world settings due to the large and dynamic website changes; (3) The model may struggle with scalability as the dataset grows.
@@
K-NN based WF Attacks: K-NN is a non-parametric, lazy learning algorithm used for classification and regression tasks. It works by finding the most similar instances (neighbors) to a given input sample and then assigning the class based on a majority vote from these neighbors. In the case of website fingerprinting, K-NN classifies network traffic based on the similarity of the features (such as packet sizes or packet ordering patterns) to the features of known traffic traces [wang2016realistically]. In 2013, Wang et al. [wang2013improved] improved the K-NN algorithm by introducing time-based splitting and classification-based splitting to isolate traffic corresponding to single or multiple web pages. Their experiments demonstrated that time-based splitting could achieve up to 97.00% TPR for detecting single-page sequences, while classification-based methods further improved accuracy in cases of overlapping traffic. Later, Wang et al. [wang2014effective] applied the K-NN classifier to a prominent feature set with weight adjustments in a realistic open-world scenario. Their results showed an 85.00% TPR with a FPR of 0.60%, outperforming previous work, which achieved 83.00% TPR with a FPR of 6.00%.
@@
Previous studies [wang2013improved, wang2014effective, wang2016realistically] often neglect hyperlink transition information, as it introduces additional “noise” that complicates the classification of the original webpage. To address this issue, Zhuo et al. [zhuo2017website] proposed a WFP technique that integrates PHMM with K-NN, where PHMM is traditionally applied in bioinformatics for DNA sequence analysis. This method accounts for hyperlink transitions, making it more realistic and effective. The technique models websites by considering different webpages and their transitions, allowing for the identification of a sequence of page loads belonging to the same website. Many WF attacks assume websites do not change over time. In reality, websites are highly dynamic, leading to significant intra-class variability that reduces classification accuracy. Cui et al. [cui2019revisiting] revisited key assumptions underlying WF attacks, which allow adversaries to infer users’ web activity despite encryption. By incorporating Markov models to capture temporal and probabilistic properties of web traffic, the paper highlights that WF attacks can remain effective under more realistic conditions. To distinguish between single-page and two-page traces, the authors employed a K-NN binary classifier. Their results show that the method can identify whether a trace contains one or two pages with an accuracy of 97.00%. Traditional K-NN classifiers typically rely on the labels of the nearest samples for classification, but the diversity of multi-tab traffic can lead to sample drift, degrading performance. To address this challenge, Zhao et al. [zhao2024towards] integrated proxy-based K-NN and sample-based K-NN to enhance multi-tab webpage identification. Their experiments showed that this hybrid model improved the multi-label recall metric by 88.60% compared to the state-of-the-art attacks.
@@
Tree-based WF Attacks: Some researchers have explored Tree-based models to improve the accuracy of WF attacks. Tree-based models use decision trees [kwon2015circuit] or ensembles of trees [xu2018multi, yin2021automated] to learn from features extracted from network traffic. These models recursively split the data based on feature thresholds, aiming to maximize the separation of traffic associated with different websites. In 2015, Kwon et al. [kwon2015circuit] used C4.5 and CART decision trees to classify Tor hidden service circuits. The authors demonstrate that even without compromising the Tor network or engaging in active traffic manipulation, it is possible to infer which hidden service a Tor client is accessing. By leveraging decision trees and WF, adversaries can deanonymize users and services with high accuracy (>>99.00%). To address the challenge of a large amount of noisy data, Hayes et al. [hayes2016k] introduced k-fingerprinting, which leverages random decision forests to classify encrypted or anonymized web traffic. The method transforms packet sequences into distinctive ”fingerprints” by mapping them to the leaf node identifiers produced by a trained random forest. It achieves impressive results, with TPR of up to 85% and FPR as low as 0.02% when identifying monitored hidden services.
@@
The accuracy of WF attack models drops drastically when tested on captured traffic traces of websites due to concept drift. To address the challenges posed by concept drift and the high cost of retraining, Attarian et al. [attarian2019adawfpa] introduced AdaWFPA, an adaptive online WF attack that overcomes the limitations of static models, whose accuracy degrades over time as website traffic patterns change. AdaWFPA leverages adaptive stream mining algorithms, Adaptive Hoeffding Tree (AHT), and Adaptive Hoeffding Option Tree (AHOT), to dynamically update its classifier in real-time, ensuring robust performance against evolving website characteristics. The method is evaluated on Tor traffic datasets, demonstrating superior accuracy, precision, and recall compared to state-of-the-art static attacks, even when tested against defenses like Walkie_Talkie [wang2017walkie].